# AURA: Autonomous Upskilling with Retrieval-Augmented Agents

### v2

@@ Abstract
Designing reinforcement learning curricula for agile robots traditionally requires extensive manual tuning of reward functions, environment randomizations, and training configurations. We introduce AURA (Autonomous Upskilling with Retrieval-Augmented Agents), a schema-validated curriculum reinforcement learning (RL) framework that leverages Large Language Models (LLMs) as autonomous designers of multi-stage curricula. AURA transforms user prompts into YAML workflows that encode full reward functions, domain randomization strategies, and training configurations. All files are statically validated before any GPU time is used, ensuring efficient and reliable execution. A retrieval-augmented feedback loop allows specialized LLM agents to design, execute, and refine curriculum stages based on prior training results stored in a vector database, enabling continual improvement over time. Quantitative experiments show that AURA consistently outperforms LLM-guided baselines in generation success rate, humanoid locomotion, and manipulation tasks. Ablation studies highlight the importance of schema validation and retrieval for curriculum quality. AURA successfully trains end-to-end policies directly from user prompts and deploys them zero-shot on a custom humanoid robot in multiple environments —capabilities that did not exist previously with manually designed controllers. By abstracting the complexity of curriculum design, AURA enables scalable and adaptive policy learning pipelines that would be complex to construct by hand.
@@ I Introduction
Curriculum reinforcement learning (RL) [1, 2, 3, 4] enables robots to master complex skills by decomposing tasks into progressively harder subtasks [5]. This approach has proven to be effective in domains like agile locomotion [6, 7, 8, 9], where single-stage learning can struggle with sparse rewards or high-dimensional exploration spaces. By splitting the learning process, agents are guided through increasingly challenging environments, improving sample efficiency and convergence.
However, designing effective multi-stage curricula remains a significant bottleneck [10, 11]. Multi-stage curricula require coordinated changes across reward shaping, randomization, simulation fidelity, and optimization; as the number of stages increase, designing transitions, verifying stability, and tuning parameters scales combinatorially (the “curse of dimensionality” [12]) and becomes brittle to human error and heuristic bias [13, 14]. Moreover, failures at any stage often derail the entire training process, making automation critical for scaling curriculum RL.
Large language models (LLMs) offer a compelling alternative to optimization-based approaches [15]. Their high-level reasoning and generalization capabilities have demonstrated success in robotic planning, perception, and code generation [16, 17, 18, 19, 20, 21]. While LLMs are not directly deployable on real-time control loops due to latency constraints [22], LLM’s have shown potential in designing training schemes for real-time robot control policies [23, 24, 25]. However, current LLM-based RL pipelines inefficiently use computation on launching parallel environments to deal with malformed generations and fail to learn from experience across tasks.
What is missing is a principled framework for transforming high-level prompts into reliable, executable RL training pipelines that improve with experience. Such a system should (1) ensure syntactic and semantic validity before execution, (2) use past experience to optimize future performance, and (3) modularize the pipeline into verifiable components [26]. Without these capabilities, LLMs are limited to more fragile, trial-and-error reward design rather than scalable curriculum generation.
@@
Zero-shot deployment: Experiments showing AURA’s policies transfer zero-shot to a custom kid-sized humanoid, Berkeley Humanoid, Booster T1, [27], and manipulators, showing end-to-end prompt-to-policy ability.
@@ II-A Large Language Models in Robotics
Early efforts to link natural language to robot control and task-planning frameworks translate natural-language commands into symbolic action graphs, as demonstrated by SayCan and Code-as-Policies [17, 18]. Recent Vision–Language–Action paradigms, such as RT-1/2, , Octo, and OTTER, [16, 28, 29, 30, 31], demonstrate that pretraining on large-scale robot data yields policies that generalize across many robotic skills. Further work has extended this to cross-domain generalization, deformable object manipulation, and long-horizon household tasks, exemplified by Gemini Robotics and embodied LLMs [21, 22]. Other lines of work have explored LLM-driven planning and language-augmented reasoning for robotic control pipelines [19, 20, 21, 32, 33, 34, 35]. Despite their semantic flexibility, these methods continue to rely on hand-tuned, low-level controllers and standard actuation assumptions, with LLMs typically external to the closed control loop [32, 33]. RoboGen addresses these limitations through decomposing long-horizon tasks into skills, which are tuned or trained by LLMs [36].
@@ II-B Curriculum Reinforcement Learning
Reinforcement learning often struggles with sparse rewards and long‐horizon problems, where meaningful feedback is rare or delayed over many timesteps [37]. Curriculum RL addresses sparse-reward and long-horizon problems by exposing agents to a progression of gradually more difficult tasks [1, 5, 14]. Surveys and benchmarking studies note that practical curriculum design remains heuristic and sensitive to context-specific particulars [14]. Reverse-curriculum approaches [38, 39] employ a staged training process that progressively increases the task’s difficulty from previous successful demonstrations, thereby enhancing exploration and sample efficiency. Domain-randomization and curriculum-based RL have enabled real-world legged locomotion and manipulation in challenging environments [2, 3, 4, 6, 7, 8, 9, 40, 41, 42, 43, 44].
@@ II-C LLM-guided RL and Curricula
Emerging systems such as CurricuLLM [24] and Eureka [23, 45] utilize LLMs to generate reward functions for RL training. Eureka uses evolutionary search and prompts LLMs in parallel to directly generate reward code for various tasks in simulation. DrEureka extends Eureka to generate domain randomizations and consider hardware safety, demonstrating real-world locomotion tasks on a Unitree Go1 quadruped robot [25]. CurricuLLM [24] similarly uses evolutionary search and directly generates code, but focuses on curriculum generation and humanoid locomotion, showing real-world success on the Berkeley Humanoid [27].
These frameworks require sampling a large number of training runs to obtain a viable or high-performing policy, relying on LLM stochasticity and brute-force search. RAG and VDBs are approaches to enhance the LLM knowledge beyond what the LLM has been trained on [46], which AURA utilizes to construct effective curricula.
@@ III-A Problem Setup and Curriculum Formalism
We define a curriculum as an ordered sequence of training stages, , where each stage is comprised of a reward , domain-randomization distributions over environment parameters , a training configuration , and a promotion criterion that governs advancement. The curriculum stages are modeled as an MDP , [47, 48] with denoting randomized physics properties [49, 50]. AURA generates for each stage as schema-compliant YAML files that are statically validated and compiled into an MJX-based RL pipeline [51, 52]. We use this notation to describe generation, validation, training, and iteration.
@@ III-C Retrieval-Augmented High-Level Planning
AURA begins by leveraging LLMs and a vector database of prior curricula and outcomes to generate a high-level plan. The user-provided task description and robot specification are first passed into a query LLM, which synthesizes them into a structured retrieval query. This query is then encoded with OpenAI’s text-embedding-3-large [53] and used to perform cosine-similarity search over the VDB (Pinecone). From the top- most relevant prior curricula, the database returns reward, randomization, and training files, along with user feedback. A selector LLM then chooses the single most relevant example from this set, based on both user feedback and curricula relevance to the current task; this example is included in the high-level planner’s context for generating the new curricula.
@@ IV-A1 Custom Humanoid
The custom humanoid locomotion environment exposes the velocity command, joint positions, projected gravity vector, last action, phase, and max foot height as observations and actuator gains, friction coefficients, body mass, COM position, foot contact geometry, initialization configuration, and roughness as domain randomization parameters. The simulation setup was adapted from the model in [54]. AURA is run for five iterations, with each iteration generating a new curriculum. Custom humanoid locomotion is evaluated with AURA Tune, which starts with the Berkeley Humanoid reward and adapts it into a curriculum that works on the custom humanoid hardware, and AURA Blind, which generates a curriculum from scratch. All environments are run at Hz, and the curriculum is limited to M training timesteps.
@@ IV-A2 Berkeley Humanoid
For the Berkeley humanoid task, AURA uses the MuJoCo Playground’s Berkeley Humanoid environment. Berkeley Humanoid’s MuJoCo Playground policy is trained as detailed in [55]. The environment is run at 50Hz and both policies train over M steps.
@@ IV-A3 BoosterT1 Humanoid
For BoosterT1 locomotion, AURA is trained and evaluated in the MuJoCo Playground BoosterT1 simulation environment, is run at 50Hz, and trains for M steps. BoosterT1’s MuJoCo Playground policy is trained as detailed in [55].
@@ IV-D Real-World Experiments
Hardware Setup: We use a custom m-tall agile humanoid with 5-DoF legs and 3-DoF arms [56]. RL policies run at Hz with sensors and actuators at Hz. The end-to-end policy uses no estimators or low-level control beyond servo internal impedance controllers.
Fig. 1 shows a policy generated by AURA deployed on real hardware in a zero-shot setting. The outdoor locomotion policy shown in the figure is trained by AURA Tune, while the robot traversing a step is trained by AURA Blind. In response to a user prompt requesting outdoor-capable locomotion, AURA automatically generated a four-stage curriculum progressing from flat terrain with low randomization to rough terrain with high randomization, diverse velocity commands, and significant external perturbations. The resulting policy demonstrated robust zero-shot generalization when deployed outdoors, successfully handling uneven terrain and maintaining consistent gait tracking. The robot achieved walking at and average gait frequency of . It withstood substantial perturbations, recovering from lateral pushes up to and from being physically pushed off a high platform without falling, which surpassed the performance of existing manually tuned controllers [56].
@@ V Limitations
Human Evaluation: Although AURA’s training pipeline generates curricula autonomously, it still relies on human feedback at the end of each iteration to provide qualitative assessment of the deployed policy. While naively applying Vision-Language Models often fails to accurately describe robot data [57, 58], recent frameworks demonstrate promising methods for automatically analyzing robot trajectories and delivering feedback [59, 60]. These could be used in future work to fully automate feedback between iterations.

### v3

@@ Abstract
Designing reinforcement learning curricula for agile robots traditionally requires extensive manual tuning of reward functions, environment randomizations, and training configurations. We introduce AURA (Autonomous Upskilling with Retrieval-Augmented Agents), a schema-validated curriculum reinforcement learning (RL) framework that leverages Large Language Models (LLMs) as autonomous designers of multi-stage curricula. AURA transforms user prompts into YAML workflows that encode full reward functions, domain randomization strategies, and training configurations. All files are statically validated before any GPU time is used, ensuring efficient and reliable execution. A retrieval-augmented feedback loop allows specialized LLM agents to design, execute, and refine curriculum stages based on prior training results stored in a vector database, enabling continual improvement over time. Quantitative experiments show that AURA consistently outperforms LLM-guided baselines in generation success rate, humanoid locomotion, and manipulation tasks. Ablation studies highlight the importance of schema validation and retrieval for curriculum quality. AURA successfully trains end-to-end policies directly from user prompts and deploys them zero-shot on a custom humanoid robot in multiple environments —capabilities that did not exist previously with manually designed controllers. By abstracting the complexity of curriculum design, AURA enables scalable and adaptive policy learning pipelines that would be complex to construct by hand. aura-research.org
@@ I Introduction
Curriculum reinforcement learning (RL) [narvekar2020curriculum, Li2024RLbipedal, anymalparkourlearningagile, fourierlatentdynamics] enables robots to master complex skills by decomposing tasks into progressively harder subtasks [rl_curriculum]. This approach has proven to be effective in domains like agile locomotion [sciencesoccer, real_world_digit, massive_rl, learningtorquecontrolquadrupedal], where single-stage learning can struggle with sparse rewards or high-dimensional exploration spaces. By splitting the learning process, agents are guided through increasingly challenging environments, improving sample efficiency and convergence.
However, designing effective multi-stage curricula remains a significant bottleneck [jang2018combining, nguyen2021robots]. Multi-stage curricula require coordinated changes across reward shaping, randomization, simulation fidelity, and optimization; as the number of stages increase, designing transitions, verifying stability, and tuning parameters scales combinatorially (the “curse of dimensionality” [li2025causally]) and becomes brittle to human error and heuristic bias [sayar2024diffusionCurriculum, soviany2022curriculum]. Moreover, failures at any stage often derail the entire training process, making automation critical for scaling curriculum RL.
Large language models (LLMs) offer a compelling alternative to optimization-based approaches [hu2020learning]. Their high-level reasoning and generalization capabilities have demonstrated success in robotic planning, perception, and code generation [brohan2022rt, ahn2022can, liang2023code, sun2024prompt, mower2024ros, mon2025embodied]. While LLMs are not directly deployable on real-time control loops due to latency constraints [team2025gemini], LLM’s have shown potential in designing training schemes for real-time robot control policies [ma2023eureka, ryu2024curricullm, ma2024dreureka]. However, current LLM-based RL pipelines inefficiently use computation on launching parallel environments to deal with malformed generations and fail to learn from experience across tasks.
What is missing is a principled framework for transforming high-level prompts into reliable, executable RL training pipelines that improve with experience. Such a system should (1) ensure syntactic and semantic validity before execution, (2) use past experience to optimize future performance, and (3) modularize the pipeline into verifiable components [Erak_2024]. Without these capabilities, LLMs are limited to more fragile, trial-and-error reward design rather than scalable curriculum generation.
@@
Zero-shot deployment: Experiments showing AURA’s policies transfer zero-shot to a custom kid-sized humanoid, Berkeley Humanoid, Booster T1, [berkeleyhumanoid], and manipulators, showing end-to-end prompt-to-policy ability.
@@ II-A Large Language Models in Robotics
Early efforts to link natural language to robot control and task-planning frameworks translate natural-language commands into symbolic action graphs, as demonstrated by SayCan and Code-as-Policies [ahn2022can, liang2023code]. Recent Vision–Language–Action paradigms, such as RT-1/2, , Octo, and OTTER, [brohan2022rt, zitkovich2023rt, pi0, octo, otter], demonstrate that pretraining on large-scale robot data yields policies that generalize across many robotic skills. Further work has extended this to cross-domain generalization, deformable object manipulation, and long-horizon household tasks, exemplified by Gemini Robotics and embodied LLMs [mon2025embodied, team2025gemini]. Other lines of work have explored LLM-driven planning and language-augmented reasoning for robotic control pipelines [sun2024prompt, mower2024ros, mon2025embodied, llm_robotics_survey, ai_review_humaniod, llm_robot_task_plan, rho2024language]. Despite their semantic flexibility, these methods continue to rely on hand-tuned, low-level controllers and standard actuation assumptions, with LLMs typically external to the closed control loop [llm_robotics_survey, ai_review_humaniod]. RoboGen addresses these limitations through decomposing long-horizon tasks into skills, which are tuned or trained by LLMs [auerbach2014robogen].
@@ II-B Curriculum Reinforcement Learning
Reinforcement learning often struggles with sparse rewards and long‐horizon problems, where meaningful feedback is rare or delayed over many timesteps [discover]. Curriculum RL addresses sparse-reward and long-horizon problems by exposing agents to a progression of gradually more difficult tasks [narvekar2020curriculum, rl_curriculum, soviany2022curriculum]. Surveys and benchmarking studies note that practical curriculum design remains heuristic and sensitive to context-specific particulars [soviany2022curriculum]. Reverse-curriculum approaches [reverse_curriculum_rl, tao2024reverse] employ a staged training process that progressively increases the task’s difficulty from previous successful demonstrations, thereby enhancing exploration and sample efficiency. Domain-randomization and curriculum-based RL have enabled real-world legged locomotion and manipulation in challenging environments [Li2024RLbipedal, anymalparkourlearningagile, fourierlatentdynamics, sciencesoccer, real_world_digit, massive_rl, learningtorquecontrolquadrupedal, zhuang2024humanoidparkourlearning, quadsoccershoots, dribblebot, nimbro_2022, adu2023exploring].
@@ II-C LLM-guided RL and Curricula
Emerging systems such as CurricuLLM [ryu2024curricullm] and Eureka [ma2023eureka, liang2024eurekaverse] utilize LLMs to generate reward functions for RL training. Eureka uses evolutionary search and prompts LLMs in parallel to directly generate reward code for various tasks in simulation. DrEureka extends Eureka to generate domain randomizations and consider hardware safety, demonstrating real-world locomotion tasks on a Unitree Go1 quadruped robot [ma2024dreureka]. CurricuLLM [ryu2024curricullm] similarly uses evolutionary search and directly generates code, but focuses on curriculum generation and humanoid locomotion, showing real-world success on the Berkeley Humanoid [berkeleyhumanoid].
These frameworks require sampling a large number of training runs to obtain a viable or high-performing policy, relying on LLM stochasticity and brute-force search. RAG and VDBs are approaches to enhance the LLM knowledge beyond what the LLM has been trained on [liu2025hm], which AURA utilizes to construct effective curricula.
@@ III-A Problem Setup and Curriculum Formalism
We define a curriculum as an ordered sequence of training stages, , where each stage is comprised of a reward , domain-randomization distributions over environment parameters , a training configuration , and a promotion criterion that governs advancement. The curriculum stages are modeled as an MDP , [sutton1998reinforcement, puterman2014markov] with denoting randomized physics properties [tobin2017domainrandomization, muratore2021data]. AURA generates for each stage as schema-compliant YAML files that are statically validated and compiled into an MJX-based RL pipeline [todorov2012mujoco, brax2021github]. We use this notation to describe generation, validation, training, and iteration.
@@ III-C Retrieval-Augmented High-Level Planning
AURA begins by leveraging LLMs and a vector database of prior curricula and outcomes to generate a high-level plan. The user-provided task description and robot specification are first passed into a query LLM, which synthesizes them into a structured retrieval query. This query is then encoded with OpenAI’s text-embedding-3-large [openai_text_embedding_v3] and used to perform cosine-similarity search over the VDB (Pinecone). From the top- most relevant prior curricula, the database returns reward, randomization, and training files, along with user feedback. A selector LLM then chooses the single most relevant example from this set, based on both user feedback and curricula relevance to the current task; this example is included in the high-level planner’s context for generating the new curricula.
@@ IV-A1 Custom Humanoid
The custom humanoid locomotion environment exposes the velocity command, joint positions, projected gravity vector, last action, phase, and max foot height as observations and actuator gains, friction coefficients, body mass, COM position, foot contact geometry, initialization configuration, and roughness as domain randomization parameters. The simulation setup was adapted from the model in [bruce_parallel_sim]. AURA is run for five iterations, with each iteration generating a new curriculum. Custom humanoid locomotion is evaluated with AURA Tune, which starts with the Berkeley Humanoid reward and adapts it into a curriculum that works on the custom humanoid hardware, and AURA Blind, which generates a curriculum from scratch. All environments are run at Hz, and the curriculum is limited to M training timesteps.
@@ IV-A2 Berkeley Humanoid
For the Berkeley humanoid task, AURA uses the MuJoCo Playground’s Berkeley Humanoid environment. Berkeley Humanoid’s MuJoCo Playground policy is trained as detailed in [mujoco_playground_2025]. The environment is run at 50Hz and both policies train over M steps.
@@ IV-A3 BoosterT1 Humanoid
For BoosterT1 locomotion, AURA is trained and evaluated in the MuJoCo Playground BoosterT1 simulation environment, is run at 50Hz, and trains for M steps. BoosterT1’s MuJoCo Playground policy is trained as detailed in [mujoco_playground_2025].
@@ IV-D Real-World Experiments
Hardware Setup: We use a custom m-tall agile humanoid with 5-DoF legs and 3-DoF arms [bruce]. RL policies run at Hz with sensors and actuators at Hz. The end-to-end policy uses no estimators or low-level control beyond servo internal impedance controllers.
Fig. 1 shows a policy generated by AURA deployed on real hardware in a zero-shot setting. The outdoor locomotion policy shown in the figure is trained by AURA Tune, while the robot traversing a step is trained by AURA Blind. In response to a user prompt requesting outdoor-capable locomotion, AURA automatically generated a four-stage curriculum progressing from flat terrain with low randomization to rough terrain with high randomization, diverse velocity commands, and significant external perturbations. The resulting policy demonstrated robust zero-shot generalization when deployed outdoors, successfully handling uneven terrain and maintaining consistent gait tracking. The robot achieved walking at and average gait frequency of . It withstood substantial perturbations, recovering from lateral pushes up to and from being physically pushed off a high platform without falling, which surpassed the performance of existing manually tuned controllers [bruce].
@@ V Limitations
Human Evaluation: Although AURA’s training pipeline generates curricula autonomously, it still relies on human feedback at the end of each iteration to provide qualitative assessment of the deployed policy. While naively applying Vision-Language Models often fails to accurately describe robot data [robo2vlm, robovqa], recent frameworks demonstrate promising methods for automatically analyzing robot trajectories and delivering feedback [autoeval, rewind]. These could be used in future work to fully automate feedback between iterations.
@@ VI Conclusion
-A Implementation and Testing Details
-A1 Training via Proximal Policy Optimization (PPO)
We train the control policy with PPO, which constrains the policy update to prevent destructive parameter jumps. Let
denote the probability ratio between the new and old policies, and let be the generalized advantage estimate at timestep . The clipped surrogate objective is
where is the clipping parameter (we use unless otherwise noted). During training, we maximize with Adam, using entropy regularization for exploration and a value–function loss with coefficient .
-A2 Score Calculation
During one evaluation run, we launch parallel environments and roll each for a horizon of simulation steps (or until failure). All three metrics are normalized to for direct comparability.
Survival Score.
Let denote the number of steps survived by environment . The Survival Score is the mean fractional episode length
Linear‑Velocity Tracking Score.
For step in environment let be the commanded planar velocity and the robot’s actual planar COM velocity expressed in the local frame. Define the squared tracking error
The YAML exponential_decay entry with corresponds to the per‑step reward
Aggregating over time and averaging across the batch yields the normalized Linear‑Velocity Tracking Score
Summary.
The triplet captures robustness (survival), command‑following fidelity, and gait coordination, respectively, and forms the basis of all quantitative comparisons in the main paper.
-A3 AURA Task Prompts for Training Launch
These five user prompts were used as task inputs to AURA for evaluating its curriculum training-launch success rate across the custom humanoid and Berkeley Humanoid embodiments. For baseline comparisons, we used the prompts provided in each baseline’s open-source repository to generate their respective training files.
-A4 Training-Launch Comparison Details
In the following section, we provide a more detailed description of the environments used to evaluate the CurricuLLM and Eureka training-launch success rates. All environments were evaluated over training launches.
CurricuLLM
CurricuLLM’s training-launch success rate was evaluated on their Berkeley Humanoid environment from [ryu2024curricullm], with all environment scripts and tasks taken directly from their open-source repository.
Eureka
Eureka’s training-launch success rate was evaluated on their only legged robot environment, ANYmal, across training-launch attempts, successfully launching training times. We also tested training-launch success across all embodiments provided in their open-source repository, which include the Gymnasium environments Shadow Hand, Franka Emika Panda, Ant, Humanoid, and Cartpole, as well as other robot environments such as ANYmal, Allegro, Ball Balance, and Quadcopter.
-B AURA LLM Agent Prompt Templates
The prompts below define the instruction templates used by AURA’s LLM agents throughout the curriculum generation and training process. These templates structure how input data, such as user task prompts, stage descriptions, and past training artifacts, are transformed into schema-compliant outputs like workflows, configuration files, and reward functions. Placeholders in the form of <INSERT_...> denote dynamic inputs that are automatically populated by the AURA framework at runtime with the appropriate content, ensuring that each prompt remains generalizable while retaining full contextual relevance for the target task.
-B1 Curriculum LLM
-B2 RAG
These prompts are used within the RAG Block to query, retrieve, and select suitable past workflows and training files to be used as context for the current task.
VDB Query LLM
Selector LLM
-B3 Stage-Level LLM
-B4 Feedback LLM
The above feedback prompt, together with the next stage’s description and all the stage’s metrics log file, is prepended to the Per-Stage LLM prompt for use by the Feedback LLM.
-C AURA Outputs
Configuration YAML
Reward YAML
Randomization YAML
Reward YAML
Randomization YAML